# Self-consistent electrostatic formalism of bulk electrolytes based on the asymmetric treatment of the short- and long-range ion interactions

### v1

@@ III.3 Overscreening and underscreening
In Figs. 5(a)-(b), we display the screening parameter of monovalent liquids at two different ion sizes. One sees that the CCDH formalism can reproduce the overscreening effect ($\kappa>\kappa_{0}$) only perturbatively. However, despite being based on the leading-order perturbative solution of Eq. (61), the SCDH-I prediction (62) can accurately account for the substantial increment of the dimensionless screening parameter by salt addition up to $n_{i}\sim 0.5$ M for $d=4.6$, and up to $n_{i}\sim 0.8$ M for $d=4.0$. Beyond these concentrations, the agreement of the SCDH-I prediction with the HNC and MC data is only qualitative.
@@
Figure 5: (Color online) (a)-(c) Dimensionless screening parameter against concentration at the HC sizes and liquid compositions indicated in the legends. Solid curves: SCDH-I prediction in Eq. (62). Dashed curves: CCDH result. Disks: HNC data from (a) Fig.1 of Ref. AttardRev ($T=300$ K and $\varepsilon_{\rm w}=78.5$) and (b)-(c) Figs.10-11 of Ref. Kj2020 ($T=298$ K and $\varepsilon_{\rm w}=78.4$). Blue squares: MC data of Ref. Kj2020 . (d) Experimental data of Ref. exp (symbols) and SCDH-I prediction (curves) for the divalent ${\rm MgSO}_{4}$ and ${\rm CuSO}_{4}$ solutions ($d=3.2$), and the trivalent ${\rm LaFeCN}_{6}$ electrolyte ($d=6.3$) at the temperature $T=298$ K and dielectric permittivity $\varepsilon_{\rm w}=78.4$.
Fig. 5(c) displays the reverse case of divalent solutions characterized by strong counterion attraction favoring the formation of ion pairs. The pair formation partially neutralizing the liquid lowers its charge strength and reduces its screening ability Kj1 ; Kj2 ; Kj1995 ; Kj2020 . In agreement with HNC data, this underscreening effect ($\kappa<\kappa_{0}$) is reproduced by the CCDH and SCDH-I formalisms up to $n_{i}\sim 10$ mM. However, the prediction of these three approaches deviate from the MC data at lower concentrations. As a result, the reversal of the underscreening trend displayed by the MC data is significantly exaggerated by our formalisms as well as by the HNC approach.
Finally, in Fig. 5(d), we confront the screening parameter (62) of the SCDH-I formalism with the experimental data of Ref. exp for the divalent ${\rm MgSO}_{4}$ and ${\rm CuSO}_{4}$ solutions, and the trivalent ${\rm LaFeCN}_{6}$ electrolyte. The figure shows that by using the hydrated ion size as the only fitting parameter (see the caption), our approach can reproduce with reasonable accuracy the underscreening effect displayed by the experimental data for dilute salt. At larger concentrations, the departure from the experimental data accompanied with the exaggerated reversal of the curve trend is similar to that observed in Fig. 5(c). The access to the corresponding strong salt regime requires the use of calculation schemes beyond the first-order cumulant closure presented in Sec. II.6. The upgraded versions of our formalism using higher order closures will be presented in upcoming works.
@@ IV Conclusions
For monovalent electrolytes with typical hydrated ion sizes $d\sim 4$, the asymmetric treatment of the short- and long-range ion interactions allows the accurate prediction of the HC-dominated osmotic pressure and the electrostatic correlation-driven internal energy up to the molar concentration regime. This is a substantial upgrade of our earlier CCDH formalism. In the case of multivalent ions, the quantitative accuracy of the SCDH-I formalism is limited to submolar concentrations. However, even in the vicinity of the molar concentration range, the SCDH-I approach improves significantly the qualitative agreement of the CCDH formalism with MC simulations.
Within the framework of our functional integral theory, we reproduced as well the characteristic Eq. (62) of the screening parameter previously derived by the dressed-ion formalism Kj1995 . The leading-order perturbative solution of this equation has been compared with HNC and MC results as well as with experimental data. We found that for monovalent ions with typical sizes, the SCDH-I approach can accurately reproduce the overscreening effect close to the molar concentration regime. In the case of multivalent electrolytes, the accuracy of our formalism in predicting the underlying underscreening effect is limited to strictly dilute salt concentrations. With the aim to extend our results to strong salt conditions, we are currently developing upgraded versions of our formalism using high order closures of the SD Eq. (28). These improvements will be reported in future works.

### v2

@@ III.4 Overscreening and underscreening
Figure 7: (Color online) (a)-(c) Dimensionless screening parameter against concentration at the HC sizes and liquid compositions indicated in the legends. Solid curves: SCDH prediction in Eq. (62). Dashed curves: CCDH result. Disks: HNC data from (a) Fig.1 of Ref. AttardRev ($T=300$ K and $\varepsilon_{\rm w}=78.5$) and (b)-(c) Figs.10-11 of Ref. Kj2020 ($T=298$ K and $\varepsilon_{\rm w}=78.4$). Blue squares: MC data of Ref. Kj2020 . (d) Experimental data of Ref. exp (symbols) and SCDH prediction (curves) for the divalent ${\rm MgSO}_{4}$ and ${\rm CuSO}_{4}$ solutions ($d=3.2$), and the trivalent ${\rm LaFeCN}_{6}$ electrolyte ($d=6.3$) at the temperature $T=298$ K and dielectric permittivity $\varepsilon_{\rm w}=78.4$.
In Figs. 7(a)-(b), we display the screening parameter of monovalent liquids at two different ion sizes. One sees that the CCDH formalism can reproduce the overscreening effect ($\kappa>\kappa_{0}$) only perturbatively. However, despite being based on the leading-order perturbative solution of Eq. (61), the SCDH prediction (62) can accurately account for the substantial increment of the dimensionless screening parameter by salt addition up to $n_{i}\sim 0.5$ M for $d=4.6$, and up to $n_{i}\sim 0.8$ M for $d=4.0$. Beyond these concentrations, the agreement of the SCDH prediction with the HNC and MC data is only qualitative.
@@
Fig. 7(c) displays the reverse case of divalent solutions characterized by strong counterion attraction favoring the formation of ion pairs. The pair formation partially neutralizing the liquid lowers its charge strength and reduces its screening ability Kj1 ; Kj2 ; Kj1995 ; Kj2020 . In agreement with HNC data, this underscreening effect ($\kappa<\kappa_{0}$) is reproduced by the CCDH and SCDH formalisms up to $n_{i}\sim 10$ mM. However, the prediction of these three approaches deviate from the MC data at lower concentrations. As a result, the reversal of the underscreening trend displayed by the MC data is significantly exaggerated by our formalisms as well as by the HNC approach.
Finally, in Fig. 7(d), we confront the screening parameter (62) of the SCDH formalism with the experimental data of Ref. exp for the divalent ${\rm MgSO}_{4}$ and ${\rm CuSO}_{4}$ solutions, and the trivalent ${\rm LaFeCN}_{6}$ electrolyte. The figure shows that by using the hydrated ion size as the only fitting parameter (see the caption), our approach can reproduce with reasonable accuracy the underscreening effect displayed by the experimental data for dilute salt. At larger concentrations, the departure from the experimental data accompanied with the exaggerated reversal of the curve trend is similar to that observed in Fig. 7(c). The access to the corresponding strong salt regime requires the use of calculation schemes beyond the first-order cumulant closure presented in Sec. II.6.
At this point, we note that the effect of salt-induced dielectric decrement BuyukDielDec may also contribute to the discrepancy between the theoretical curves and the experimental data in the large concentration regime of Fig. 7(d). In future works, the corresponding feature may be taken into account by using salt-dependent dielectric permittivities NetzMC2 ; Buyuk2020 or via the incorporation of explicit solvent into the present formalism BuyukDielDec .
@@ IV Conclusions
For monovalent electrolytes with typical hydrated ion sizes $d\sim 4$, the asymmetric treatment of the short- and long-range ion interactions allows the accurate prediction of the HC-dominated osmotic pressures, the electrostatic correlation-driven internal energies, and the ionic pair distributions governed by the competition between electrostatic and HC interactions up to the molar concentration regime. This is a substantial upgrade of our earlier CCDH formalism. In the case of multivalent ions, the quantitative accuracy of the SCDH formalism is limited to submolar concentrations. However, even in the vicinity of the molar concentration range, the SCDH approach improves significantly the qualitative agreement of the CCDH formalism with MC simulations.
Within the framework of our functional integral theory, we reproduced as well the characteristic Eq. (62) of the screening parameter previously derived by the dressed-ion formalism Kj1995 . The leading-order perturbative solution of this equation has been compared with HNC and MC results as well as with experimental data. We found that for monovalent ions with typical sizes, the SCDH approach can accurately reproduce the overscreening effect close to the molar concentration regime. In the case of multivalent electrolytes, the accuracy of our formalism in predicting the underlying underscreening effect is limited to strictly dilute salt concentrations. With the aim to extend our results to strong salt conditions, we are currently developing upgraded versions of our formalism based on high order closures of the SD Eq. (28) and incorporating explicit solvent BuyukDielDec . These improvements will be reported in future works.
Data availability
The Author declares that the parameters and data supporting the graphically illustrated results are available within the article. The bibliographic references of the HNC, MC, and experimental data taken from the literature are cited in the caption of each figure.
Conflict of interest
The author declares no competing interests.